# On the geometry of the free factor graph for ${\rm{Aut}}(F_{N})$

### 06 Dec 2023 v1

@@ 5 Problem list
Our results extend the list of automorphisms that are known to have undistorted orbits in $\mathcal{AF}_{N}$: Bestvina and Feighn [BF14] proved this is true for fully irredicible automorphisms, and we have added inner automorphisms ${\rm{ad}}_{b}$ given by filling elements.
@@
The quasi-flats in $\mathcal{AF}_{N}$ constructed in this paper come in families where each pair of quasi-flats coarsely share a common line: if $b$ is the boundary word for a surface $\Sigma$, then independent psuedo-Anosov homeomorphisms of $\Sigma$ will give distinct quasi-flats, each pair of which will have bounded neighbourhoods whose intersection is a quasi-line that is a bounded neighbourhood of an $\langle{\rm{ad}}_{b}\rangle$-orbit. This observation allows us to construct many more non-equivariant quasi-isometric embeddings $\mathbb{Z}^{2}\hookrightarrow\mathcal{AF}_{N}$ by piecing together half-flats and parallel strips from the various equivariant flats coarsely containing $\langle{\rm{ad}}_{b}\rangle$-orbits. This pattern of intersections shows that the families of quasi-flats that we currently have cannot be used to define a relative hyperbolic structure on $\mathcal{AF}_{N}$. (We refer to [Sis12] for the definition of such a structure.) But our lack of knowledge about quasi-flats in general means that we cannot answer the following question.

### 12 Mar 2024 v2

@@ 5 Problem list
Our results extend the list of automorphisms that are known to have undistorted orbits in $\mathcal{AF}_{N}$: Bestvina and Feighn [BF14] proved this is true for fully irreducible automorphisms, and we have added inner automorphisms ${\rm{ad}}_{b}$ given by filling elements.
@@
The quasi-flats in $\mathcal{AF}_{N}$ constructed in this paper come in families where each pair of quasi-flats coarsely share a common line: if $b$ is the boundary word for a surface $\Sigma$, then independent pseudo-Anosov homeomorphisms of $\Sigma$ will give distinct quasi-flats, each pair of which will have bounded neighbourhoods whose intersection is a quasi-line that is a bounded neighbourhood of an $\langle{\rm{ad}}_{b}\rangle$-orbit. This observation allows us to construct many more non-equivariant quasi-isometric embeddings $\mathbb{Z}^{2}\hookrightarrow\mathcal{AF}_{N}$ by piecing together half-flats and parallel strips from the various equivariant flats coarsely containing $\langle{\rm{ad}}_{b}\rangle$-orbits. This pattern of intersections shows that the families of quasi-flats that we currently have cannot be used to define a relative hyperbolic structure on $\mathcal{AF}_{N}$. (We refer to [Sis12] for the definition of such a structure.) But our lack of knowledge about quasi-flats in general means that we cannot answer the following question.